At what rate must motorist A travel to overtake motorist B who is traveling at a rate 20 miles/hour slower if A starts two hours after B and wishes to overtake B in 4 hours?
Answer by ankor@dixie-net.com(22740)

At what rate must motorist A travel to overtake motorist B who is traveling at a rate 20 miles/hour slower if A starts two hours after B and wishes to overtake B in 4 hours?
:
let r = the rate required by A to overtake B
Then
(r-20) = the rate of B
:
Travel time of A is 4 hrs, travel time of B is 6 hrs
:
Write a distance equation, both will have traveled the same dist when A catches B
dist = speed*time
:
4r = 6(r-20)
4r = 6r - 120
120 = 6r - 4r
r = 120/2
r = 60 mph, to overtake in 4 hrs
:
:
Confirm this by finding the distance each traveled, should be equal
4(60) = 240 mi
6(40) = 240 mi